Answer problem #2, showing ALL work accordingly. Thank you.
Problem 2 Use the sum-difference formulas to find sin(75º), cos(75°), and tan(75)
Answer #1

we have the formulas as

\sin(x+y)= \sin(x)\cos(y)+\cos(x)\sin(y)

\cos(x+y)= \cos(x)\cos(y)-\sin(x)\sin(y)

\tan(x+y)= \frac{\sin(x+y)}{\cos(x+y)}

and also we have the values,

\sin(45\degree)= \cos(45\degree)= \frac{1}{\sqrt{2}}

\sin(30\degree)= \frac{1}{2}

\cos(30\degree)= \frac{\sqrt{3}}{2}

so we get the values as,

\sin(75\degree)= \sin(45\degree+30\degree)= \sin(45\degree)\cos(30\degree)+\cos(45\degree)\sin(30\degree)

= (\frac{1}{\sqrt{2}}*\frac{\sqrt{3}}{2})+(\frac{1}{\sqrt{2}}*\frac{1}{2})

=\boldsymbol{ \frac{\sqrt{3}+1}{2\sqrt{2}}}*\frac{\sqrt{2}}{\sqrt{2}}

  = \boldsymbol{\frac{\sqrt{6}+\sqrt{2}}{4}}

\cos(75\degree)= \cos(45\degree+30\degree)= \cos(45\degree)\cos(30\degree)-\sin(45\degree)\sin(30\degree)

=( \frac{1}{\sqrt{2}}*\frac{\sqrt{3}}{2})-(\frac{1}{\sqrt{2}}*\frac{1}{2})

  = \boldsymbol{\frac{\sqrt{3}-1}{2\sqrt{2}}}*\frac{\sqrt{2}}{\sqrt{2}}

= \boldsymbol{\frac{\sqrt{6}-\sqrt{2}}{4}}

\tan(75\degree)= \frac{\sin(75\degree)}{\cos(75\degree)}= \frac{\sqrt{3}+1}{\sqrt{3}-1}

  = \frac{\sqrt{3}+1(\sqrt{3}+1)}{\sqrt{3}-1(\sqrt{3}+1)}= \frac{4+2\sqrt{3}}{3-1}

=\boldsymbol{ 2+\sqrt{3}}